OpenStudy (3psilon):

Two vectors A and B have magnitudes of 6m and make an angle of 60 degrees with each other. Find A DOT B.

OpenStudy (anonymous):

\[\cos \theta=\frac{ x \dot y }{ \left| x \right|\left| y \right|}\]

OpenStudy (anonymous):

so then just plug in your info, and solve for x dot y, or a dot b in this case.

OpenStudy (3psilon):

How did you get the Cos theta = that stuff?

OpenStudy (anonymous):

its just a theorem or an equation which comes from the law of cosines

OpenStudy (3psilon):

So is it just 3?

OpenStudy (anonymous):

well i took it when i first read it that A and B each had a magnitude of 6, but if they are 6 together, then yes it would be 3.

OpenStudy (3psilon):

Oh they each have 6 So it would be 18>?

OpenStudy (3psilon):

18